OpenStudy (anonymous):

Find the area of a triangle with a base of 16m and an altitude of 7m: a.15 b.28 c.56 d.112

OpenStudy (mathstudent55):

A = (1/2)bh where b = base and h = height

OpenStudy (mathstudent55):

height = altitude

OpenStudy (anonymous):

I'm betting $ 50 on @ajpricess

OpenStudy (ajprincess):

Area of a triangle=(1/2)*base*altitude. does that help? @angsmallworlds

OpenStudy (anonymous):

yeah i understand it now, thank you!

OpenStudy (ajprincess):

Welcome:)